# Higher property T and below-rank phenomena of lattices

### v1

@@ 1. Introduction
Property $(T_{1})$ is equivalent to property $T_{1}]$, and both properties are equivalent to Kazhdan’s property T. The fact that higher rank lattices222 Higher rank lattices are lattices in groups of rank $r\geq 2$. have property T is due to Kazhdan. For $F$ non-archimedean Theorem 1 is due to Dymara-Januszkiewicz [Dy-Jan] and Grinbaum-Reizis-Oppenheim [GR-Opp] (using the so called “Garland Method”). The general case is due to [badsau]. The proof will be sketched in §3.1.
@@
In view of [harder], it is highly likely that the characteristic 0 assumption on $F$ in Theorem 1 could be removed. See §3.3.1 for a further discussion.
@@
Over a archimedean field $F$, the groups $G$ and lattices $\Gamma$ as in Theorem 1 do not typically satisfy $T_{r-1}]$. For example $H^{2}(\operatorname{Sp}_{n}(\mathbb{Z}),\mathbb{C})\neq 0$.
@@
In short, we regard this conjecture as saying that $\Gamma$ has the super-reflexive property $(T_{r-1})$. The case $j=1$ (which was conjectured in [BFGM]) was proved recently in [Oppenheim] and [Sa-Le-at], see Theorem 61 below. When removing the assumption $V^{\Gamma}=0$ but leaving as is the rest of the setup of Conjecture 5, using induction techniques developed in [badsau], the conjecture implies that for every $j<r$, we have
@@
In case $G$ is a Lie group with property T and $V$ is unitary, the conjecture is proved in [badsau, Theorem D]. We expect that the proof of Theorem 1 could be also extended to the S-semisimple case, see Conjecture 78 below, which regards an extension of Theorem 77 by Leuzinger-Young to this setting. We also expect that the property T assumption in [badsau, Theorem D] could be removed, see [badsau, Conjecture 1.5]. This is due to the Strong Spectral Gap Conjecture 98 that will be discussed in §3.4.1. We stress that Conjecture 6 is open even in degree $1$. The degree $1$ vanishing over unitary representations, which is relevant for irreducible lattices in product of rank $1$ groups, is quite intriguing. It is equivalent to Shalom’s Spectral Gap Conjecture 97. Its validity implies for instance the full solution of Connes’ Character Rigidity conjecture and Lubotzky’s property $\tau$ conjecture. See §3.4.1 for details.
In the forthcoming paper [BBBS], together with Saar Bader and Shaked Bader, we show that lattices satisfy vanishing results for general isometric actions on von Neumann modules. In §3.1 we will sketch the proof of the next theorem and relate it to some further conjectures. For the notions of $\Gamma$-unitary von Neumann algebra and the noncommutative $L^{p}$-space $L^{p}(M)$ of a von Neumann algebra $M$ we refer to the beginning of §2.10. The reader who prefers a von-Neumann free statement should look at Corollary 8 below.
Theorem 7 ([BBBS]).
@@
Associated with any unitary representation $V$ of $\Gamma$ there is a corresponding Gaussian measure, which yields a probability measure preserving action of $\Gamma$ on a space $X$ such that $V$ embeds in $L^{2}(X)$ as a subrepresentation [zimmer]*Example 5.2.13 on p. 110. Therefore $H^{k}(\Gamma,L^{2}(X))=0$ implies $H^{k}(\Gamma,V)=0$. We could therefore express Theorem 1 in a similar way as Corollary 8 in the case $p=2$.
@@
Corollary 8 has some interesting applications. In particular, it implies conjectures of Gromov and Farb that will be dealt with in §3.2.1 and §3.3, respectively. In §3.2.2 we speculate on torsion growth, a speculation which we relate by the end of §3.3, after discussing expansions and waist inequalities, to the relations between expansion over the reals, over the integers and over finite fields. Finally, in §3.4 we regard cohomological vanishing in small degrees. We will focus mainly on degree 1 applications of Conjecture 6, which we already alluded to earlier, but also mention degree 2 application to the theory of stability.
@@
We provide a sum-of-squares characterization of property $(T_{n})$ in §2.9 for the Laplace-operator regarded as a matrix over the group ring. See §2.1 for a discussion of finiteness properties.
@@
Here $\Delta_{k}$ is the Laplace operator in degree $k$, and $\Delta_{0}$ is an element in the group ring which, viewed as a scalar, can be multiplied with matrices over the group ring. The corresponding result for $T_{n}]$ (Theorem 49) was (essentially) proved in [Bader-Nowak]. See the comment after Theorem 49 in §2.11).
@@ Acknowledgments
We thank Lubotzky for lots of advice, suggestions and encouragement. We thank Francesco Fournier Facio and Mikael de la Salle for answering questions (mentioned in the text). We thank Alon Dogon, David Fisher, Michael Glasner, Yuval Gorfine, Izhar Oppenheim, Yehuda Shalom and Stefan Witzel for insightful remarks over a first draft of this paper. In particular we thank Pierre Pansu and Raz Slutsky for a very thorough reading of this first draft and the many comments they made. Finally, we thank Saar Bader and Shaked Bader for the most delightful collaboration [BBBS] and their many inputs. In the initial phase of this project, the second author was supported by the DFG (Deutsche Forschungsgemeinschaft), project number 441426599. The second author would also like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Operators, Graphs, Groups where work on this paper was undertaken. This work was supported by EPSRC grant no EP/K032208/1.
@@ 2. Higher property T as an abstract property of groups
Unlike the next section which focuses on lattices in semisimple groups and mainly reviews recent results, this section contains new results (Theorems 11, 12 and 10), results that have only been stated in special cases in the literature before (Lemmas 21, 23, and 35) and results that give a fresh perspective to known facts (for example, Corollary 28 and §2.8).
For the most part, we assume a mild finiteness property on the groups under consideration. See Convention 17 below. The role of this assumption is discussed in §2.1 and §2.13.
@@
Another class of groups that has a chance to have higher property T are automorphism groups of free groups. The group $\operatorname{Aut}(F_{n})$ has property T for $n\geq 4$. This was proved by Nitsche [nitsche] for $n=4$, Kaluba-Nowak-Ozawa for $n=5$ [kaluba+nowak+ozawa] and by Kaluba-Kielak-Nowak for $n\geq 6$ [kaluba+kielak+nowak]. Further, Galatius showed that the rational cohomology of $\operatorname{Aut}(F_{n})$ stabilizes to that of the symmetric group $\Sigma_{n}$. As a corollary one obtains that $H^{k}(\operatorname{Aut}(F_{n}),{\mathbb{Q}})=0$ for $n\geq 2k+2$ [galatius]*Corollary 1.2. Finally, $\operatorname{Aut}(F_{n})$ has property $FA_{\lfloor\frac{2n}{3}\rfloor}$ by [bridson-helly]. According to Theorem 84 this would be a consequence of higher property T for Banach spaces, more precisely of property $(T_{\lfloor\frac{2n}{3}\rfloor})_{L^{1}}$.
@@ 2.1. Finiteness properties
Property T implies finite generation, but not finite presentation. An example due to Yves Cornulier is presented in [bekka+harpe+valette]*Theorem 3.4.2 on p. 172. The same example shows that property $T_{2}]$ does not imply finite presentation either, which we discuss next.
@@
Let $n\in\mathbb{N}$, and let $p$ be a prime. The group $\operatorname{Sp}_{n}(\mathbb{Z}[1/p])\ltimes\mathbb{Z}[1/p]^{2n}$ is not finitely presented by [bekka+harpe+valette]*Theorem 3.4.2 on p. 172. Similarly, one shows that the locally compact group $\operatorname{Sp}_{n}(\mathbb{Q}_{p})\ltimes\mathbb{Q}_{p}^{2n}$ is not compactly presented. In forthcoming work it will be shown that both groups have property $(T_{n-1})$. Moreover, the first group has a central extension that has $T_{2}]$ (cf. [bekka+harpe+valette]*Lemma 3.4.3 on p. 173).
@@
Consider Theorem 1 in the case that the field $F$ is of positive characteristic and $\Gamma$ is non-uniform lattice (which is automatic, unless $G$ is of type $A_{n}$). Then $\Gamma$ (presumably) has $T_{r-1}]$ (see Remark 3), and it is of type $FP_{r-1}({\mathbb{Z}})$, in particular of type $FP_{r-1}({\mathbb{Q}})$, by the rank theorem of Bux-Köhl-Witzel [bux+koehl+witzel]. However, the group $\Gamma$ is not of type $FP_{r}({\mathbb{Z}})$ by a theorem of Bux-Wortman [bux+wortman]*Theorem 1.2. Very likely, the same holds with $\mathbb{Z}$} replaced by $\mathbb{Q}$}.
Convention 17.
@@ 2.2. The role of the Laplace operators
Each $P_{n}$ is equipped with a ${\mathbb{C}}\Gamma$-basis defined over $\mathbb{Q}$}. We consider the cochain complex
@@ 2.3. Banach $\Gamma$-modules
The cohomology $H^{\mathbin{{\bullet}}}(\Gamma,V)$ has a canonical topological vector space structure, which is typically non-Hausdorff [guichardet]. If $\Gamma$ is of type $FP_{\infty}({\mathbb{Q}})$, which we assume (see Convention 17), and $P_{\mathbin{{\bullet}}}$ is any resolution as in (3), then $\hom_{{\mathbb{C}}\Gamma}(P_{\mathbin{{\bullet}}},V)$ is a cochain complex of Banach spaces whose topology induces the one on $H^{\mathbin{{\bullet}}}(\Gamma,V)$.
@@
Let $V$ be a Banach $\Gamma$-module. If $V$ has a non-trivial $n$-th (co)homological torsion, then $U$ has a non-trivial $n$-th (co)homological torsion for every closed inclusion of Banach $\Gamma$-modules $V\hookrightarrow U$ and for every surjective map of Banach $\Gamma$-modules $U\twoheadrightarrow V$.
@@
The statement is obvious for closed subspaces and it follows for quotients by dualizing. ∎
@@ 2.4. Ultrapowers
An additive functor between abelian categories with the property in Lemma 25 is called exact. However, the category $\mathsf{Ban}$ fails to be abelian since a bounded linear map with trivial kernel and trivial cokernel is not an isomorphism in general. Note that the (categorical) cokernel of $T\colon U\to V$ is $\operatorname{coker}(T:U\to V)=V/\overline{\operatorname{im}T}$. The category $\mathsf{Ban}$ is still an exact category. See [buehler] for a concise introduction to exact categories.
@@
By Lemma 25, we easily reduce to the case where $T$ is injective and with a dense, non-closed, image. By the open mapping theorem there exists a sequence of unit vectors $u_{i}\in U$ with $\lim T(u_{i})=0$. The element represented by $(u_{i})$ in $U_{\omega}$ is in $\ker(T_{\omega})$. Note that also $T^{*}\colon V^{*}\to U^{*}$ is injective and with a dense, non-closed, image. We thus can find a unit vector $\phi\in(V^{*})_{\omega}$ that is in $\ker((T^{*})_{\omega})$. The image of $\phi$ under the natural map $(V^{*})_{\omega}\to(V_{\omega})^{*}$ [^7] This map is an isometry into, which is surjective iff $V$ super-reflexive [Heinrich, Corollary 7.2]. vanishes on $\operatorname{im}(T_{\omega})$. Hence also on $\overline{\operatorname{im}(T_{\omega})}$. ∎
Let $F\colon\mathsf{Ban}\to\mathsf{Abelian}$ be the forgetful functor to abelian groups. The functor $F$ preserves monomorphisms or kernels but does not preserve epimorphisms or cokernels (see Remark 26). In particular, $F$ is not an exact functor. This is one of the reasons that the homological algebra underlying cohomology with coefficients in Banach spaces is a bit more unpleasant than in a purely algebraic context. The ultrapower functor $\Omega$ is helpful because Lemmas 25 and 27 imply the following corollary. It says that we can express the preservation of epimorphisms and cokernels by $F$ within $\mathsf{Ban}$, thus staying in the functional-analytic context.
@@
For a Banach $\Gamma$-module $V$, an almost invariant sequence is a sequence of vectors $v_{n}\in V$ such that for every $g\in\Gamma$, $\lim_{n}(1-g)v_{n}=0$. We end this subsection with the following observation, see [badsau]*Section 3.

### v2

@@ 1. Introduction
Property $(T_{1})$ is equivalent to property $T_{1}]$, and both properties are equivalent to Kazhdan’s property T. The fact that higher rank lattices222 Higher rank lattices are lattices in groups of rank $r\geq 2$. have property T is due to Kazhdan. For $F$ non-archimedean Theorem 1 is due to Dymara-Januszkiewicz [40] and Grinbaum-Reizis-Oppenheim [54] (using the so called “Garland Method”). The general case is due to the authors [15]. The proof will be sketched in §3.1.
@@
In view of [58], it is highly likely that the characteristic 0 assumption on $F$ in Theorem 1 could be removed. See §3.3.1 for a further discussion.
@@
Over an archimedean field $F$, the groups $G$ and lattices $\Gamma$ as in Theorem 1 do not typically satisfy $T_{r-1}]$. For example $H^{2}(\operatorname{Sp}_{n}(\mathbb{Z}),\mathbb{C})\neq 0$.
Despite the previous remark, the range of higher property T can exceed the rank. There are even rank $1$ groups with higher property T, as the following theorem demonstrates. Recall that isometry group of the Cayley hyperbolic plane, denoted $F_{4}^{-20}$, is a rank 1 simple Lie group.
Theorem 5 ([15]*Theorem B).
The group $F_{4}^{-20}$ and its lattices have property $T_{3}]$.
Noting that the 2nd and 3rd cohomologies for $F_{4}^{-20}$ with trivial coefficients vanish, see e.g [21]*§19.1, Theorem 5 follows for the Lie group and its uniform lattices by [15]*Theorem B. For non-uniform lattices it follows from [72]. We thank Antonio López Neumann for verifying this.
Remark 6.
Over a non-archimedean field, this is never the case, as the Steinberg representation is always cohomological in the degree of the rank.
@@
In short, we regard this conjecture as saying that $\Gamma$ has the super-reflexive property $(T_{r-1})$. The case $j=1$ (which was conjectured in [10]) was proved recently in [81] and [37], see Theorem 67 below. When removing the assumption $V^{\Gamma}=0$ but leaving as is the rest of the setup of Conjecture 7, using induction techniques developed in [15], the conjecture implies that for every $j<r$, we have
@@
The following two theorems from [15] establish Conjecture 8 in two important special cases. Theorem 10 is a generalization of Borel’s stability theorem to a wider range of degrees and from constant to unitary coefficients.
Theorem 9 ([15]*Theorem D).
If $G$ is a semisimple Lie group with property T and $V$ is a unitary $\Gamma$-representation, then Conjecture 8 holds true.
Theorem 10 ([15]*Theorem C).
If $G$ is a semisimple Lie group (not necessarily with property T) and $V$ is the restriction of a unitary $G$-representation, then Conjecture 8 holds true.
We expect that the proof of Theorem 1 could be also extended to the S-semisimple case (Conjecture 86 below), which is concerned with an extension of Theorem 85 by Leuzinger-Young to this setting. This is due to the Strong Spectral Gap Conjecture 106, which will be discussed in §3.4.1. We stress that Conjecture 8 is open even in degree $1$. The degree $1$ vanishing over unitary representations, which is relevant for irreducible lattices in product of rank $1$ groups, is quite intriguing. It is equivalent to Shalom’s Spectral Gap Conjecture 105. Its validity implies for instance the full solution of Connes’ Character Rigidity conjecture and Lubotzky’s property $\tau$ conjecture. See §3.4.1 for details.
In a paper [7], yet to be written together with Saar Bader and Shaked Bader, we show that lattices satisfy vanishing results for general isometric actions on von Neumann modules. In §3.1 we will sketch the proof of the next theorem and relate it to some further conjectures.
For the notions of $\Gamma$-unitary von Neumann algebra and the noncommutative $L^{p}$-space $L^{p}(M)$ of a von Neumann algebra $M$ we refer to the beginning of §2.10. The reader who prefers a von-Neumann free statement should look at Corollary 12 below.
Theorem 11 ([7]).
@@
Associated with any unitary representation $V$ of $\Gamma$ there is a corresponding Gaussian measure, which yields a probability measure preserving action of $\Gamma$ on a space $X$ such that $V$ embeds into $L^{2}(X)$ as a subrepresentation [100]*Example 5.2.13 on p. 110.
Remark 14 (Comparision of the paper [15] and the forthcoming paper [7]).
We clarify the relation between the original higher property T paper [15] and the announced paper [7]. In view of Remark 13 we could express Theorem 1, that is [15]*Theorem A, in a similar way as Corollary 12 in the case $p=2$. On the other hand, the generalization of Theorems B,C,D,E,F,G in [15], which includes Theorems 5, 9 and 10, to the setting of Banach representations remains unclear at the moment.
The forthcoming proofs of Theorem 11 and Corollary 12 do not subsume the proof of Theorem 1 but depend on it. The reason is that a comparatively simple step in [15] is no longer applicable for $p\neq 2$ and must be replaced by a completely new and delicate analysis, while the full technical machinery developed in [15] remains indispensable. See also Remark 70.
@@
The previous theorems and conjectures have lots of interesting applications and ramifications in relation to various rigidity and geometric phenomena below the rank. We discuss these in §3.2, §3.3 and §3.4. Let us highlight a few of them here.
In §3.2.1 we discuss generalizations of Borel’s stability theorem and theorems of Borel-Yang to a wider range of degrees and to unitary coefficients. These are based on Theorem 10 and certain adelic generalizations of it. Corollary 12 implies conjectures of Gromov and Farb that will be dealt with in §3.2.2 and §3.3, respectively. In §3.2.3 we speculate on torsion growth, a speculation which we relate by the end of §3.3, after discussing expansions and waist inequalities, to the relations between expansion over the reals, over the integers and over finite fields. In §3.4 we regard cohomological vanishing in small degrees. We discuss degree 1 applications of Conjecture 8, which we already alluded to earlier. Then we turn to degree 2 applications of Theorem 1 to the theory of stability. Finally, we briefly discuss an interesting use of Theorem 5 in §3.4.3
@@
We provide a sum-of-squares characterization of property $(T_{n})$ in §2.9 for the Laplace-operator regarded as a matrix over the group ring, generalizing the celebrated Ozawa criterion [83]. See §2.1 for a discussion of finiteness properties.
@@
Here $\Delta_{k}$ is the Laplace operator in degree $k$, and $\Delta_{0}$ is an element in the group ring which, viewed as a scalar, can be multiplied with matrices over the group ring. The corresponding result for $T_{n}]$ (Theorem 55) was (essentially) proved in [13]. See the comment after Theorem 55 in §2.11.
@@ Acknowledgments
We thank Lubotzky for lots of advice, suggestions and encouragement. We thank Francesco Fournier Facio, Mikael de la Salle and Antonio López Neumann for answering questions (mentioned in the text). We thank Alon Dogon, David Fisher, Michael Glasner, Yuval Gorfine, Izhar Oppenheim, Yehuda Shalom and Stefan Witzel for insightful remarks over a first draft of this paper. In particular we thank Pierre Pansu and Raz Slutsky for a very thorough reading of this first draft and the many comments they made. Finally, we thank Saar Bader and Shaked Bader for the most delightful collaboration [7] and their many inputs.
In the initial phase of this project, the second author was supported by the DFG (Deutsche Forschungsgemeinschaft), project number 441426599. The second author would also like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Operators, Graphs, Groups where work on this paper was undertaken. This work was supported by EPSRC grant no EP/K032208/1.
@@ 2. Higher property T as an abstract property of groups
Unlike the next section which focuses on lattices in semisimple groups and mainly reviews recent results, this section contains new results (Theorems 16, 17 and 15), results that have only been stated in special cases in the literature before (Lemmas 26, 28, and 40) and results that give a fresh perspective to known facts (for example, Corollary 33 and §2.8).
For the most part, we assume a mild finiteness property on the groups under consideration. See Convention 22 below. The role of this assumption is discussed in §2.1 and §2.13.
@@
Another class of groups that has a chance to have higher property T are automorphism groups of free groups. The group $\operatorname{Aut}(F_{n})$ has property T for $n\geq 4$. This was proved by Nitsche [79] for $n=4$, Kaluba-Nowak-Ozawa for $n=5$ [65] and by Kaluba-Kielak-Nowak for $n\geq 6$ [64]. Further, Galatius showed that the rational cohomology of $\operatorname{Aut}(F_{n})$ stabilizes to that of the symmetric group $\Sigma_{n}$. As a corollary one obtains that $H^{k}(\operatorname{Aut}(F_{n}),{\mathbb{Q}})=0$ for $n\geq 2k+2$ [52]*Corollary 1.2. Finally, $\operatorname{Aut}(F_{n})$ has property $FA_{\lfloor\frac{2n}{3}\rfloor}$ by [28]. According to Theorem 92 this would be a consequence of higher property T for Banach spaces, more precisely of property $(T_{\lfloor\frac{2n}{3}\rfloor})_{L^{1}}$.
@@ 2.1. Finiteness properties
Property T implies finite generation, but not finite presentation. An example due to Yves Cornulier is presented in [17]*Theorem 3.4.2 on p. 172. The same example shows that property $T_{2}]$ does not imply finite presentation either, which we discuss next.
@@
Let $n\in\mathbb{N}$, and let $p$ be a prime. The group $\operatorname{Sp}_{n}(\mathbb{Z}[1/p])\ltimes\mathbb{Z}[1/p]^{2n}$ is not finitely presented by [17]*Theorem 3.4.2 on p. 172. Similarly, one shows that the locally compact group $\operatorname{Sp}_{n}(\mathbb{Q}_{p})\ltimes\mathbb{Q}_{p}^{2n}$ is not compactly presented. In forthcoming work it will be shown that both groups have property $(T_{n-1})$. Moreover, the first group has a central extension that has $T_{2}]$ (cf. [17]*Lemma 3.4.3 on p. 173).
@@
Consider Theorem 1 in the case that the field $F$ is of positive characteristic and $\Gamma$ is non-uniform lattice (which is automatic, unless $G$ is of type $A_{n}$). Then $\Gamma$ (presumably) has $T_{r-1}]$ (see Remark 3), and it is of type $FP_{r-1}({\mathbb{Z}})$, in particular of type $FP_{r-1}({\mathbb{Q}})$, by the rank theorem of Bux-Köhl-Witzel [30]. However, the group $\Gamma$ is not of type $FP_{r}({\mathbb{Z}})$ by a theorem of Bux-Wortman [31]*Theorem 1.2. Very likely, the same holds with $\mathbb{Z}$} replaced by $\mathbb{Q}$}.
Convention 22.
@@ 2.2. The role of the Laplace operators
Each $P_{n}$ is equipped with a ${\mathbb{C}}\Gamma$-basis defined over $\mathbb{Q}$} and we assume as we may $P_{0}={\mathbb{C}}\Gamma$. We consider the cochain complex
@@ 2.3. Banach $\Gamma$-modules
The cohomology $H^{\mathbin{{\bullet}}}(\Gamma,V)$ has a canonical topological vector space structure, which is typically non-Hausdorff [57]. If $\Gamma$ is of type $FP_{\infty}({\mathbb{Q}})$, which we assume (see Convention 22), and $P_{\mathbin{{\bullet}}}$ is any resolution as in (3), then $\hom_{{\mathbb{C}}\Gamma}(P_{\mathbin{{\bullet}}},V)$ is a cochain complex of Banach spaces whose topology induces the one on $H^{\mathbin{{\bullet}}}(\Gamma,V)$.
@@
Let $V$ be a Banach $\Gamma$-module with $V^{\Gamma}$ finite dimensional and let $U<V$ be a closed sub-$\Gamma$-module. If $H^{1}(\Gamma,U)$ is not Hausdorff then also $H^{1}(\Gamma,V)$ is not Hausdorff.
@@
If $H^{1}(\Gamma,V)$ is Hausdorff then the image under $d^{1}$ of $P^{0}_{V}\cong V$ in $P^{1}_{V}$ is closed, and isomorphic to $V/V^{\Gamma}$. Since $V^{\Gamma}$ is finite dimensional, the image of $U$ is closed in $V/V^{\Gamma}$, and we deduce that the image under $d^{1}$ of $P^{0}_{U}\cong U$ in $P^{1}_{U}$ is closed, thus $H^{1}(\Gamma,U)$ is Hausdorff. ∎
@@ 2.4. Ultrapowers
An additive functor between abelian categories with the property in Lemma 30 is called exact. However, the category $\mathsf{Ban}$ fails to be abelian since a bounded linear map with trivial kernel and trivial cokernel is not an isomorphism in general. Note that the (categorical) cokernel of $T\colon U\to V$ is $\operatorname{coker}(T:U\to V)=V/\overline{\operatorname{im}T}$. The category $\mathsf{Ban}$ is still an exact category. See [29] for a concise introduction to exact categories.
@@
By Lemma 30, we easily reduce to the case where $T$ is injective and with a dense, non-closed, image. By the open mapping theorem there exists a sequence of unit vectors $u_{i}\in U$ with $\lim T(u_{i})=0$. The element represented by $(u_{i})$ in $U_{\omega}$ is in $\ker(T_{\omega})$. Note that also $T^{*}\colon V^{*}\to U^{*}$ is injective and with a dense, non-closed, image. We thus can find a unit vector $\phi\in(V^{*})_{\omega}$ that is in $\ker((T^{*})_{\omega})$. The image of $\phi$ under the natural map $(V^{*})_{\omega}\to(V_{\omega})^{*}$ [^7] This map is an isometry into, which is surjective iff $V$ super-reflexive [61, Corollary 7.2]. vanishes on $\operatorname{im}(T_{\omega})$. Hence also on $\overline{\operatorname{im}(T_{\omega})}$. ∎
Let $F\colon\mathsf{Ban}\to\mathsf{Abelian}$ be the forgetful functor to abelian groups. The functor $F$ preserves monomorphisms or kernels but does not preserve epimorphisms or cokernels (see Remark 31). In particular, $F$ is not an exact functor. This is one of the reasons that the homological algebra underlying cohomology with coefficients in Banach spaces is a bit more unpleasant than in a purely algebraic context. The ultrapower functor $\Omega$ is helpful because Lemmas 30 and 32 imply the following corollary. It says that we can express the preservation of epimorphisms and cokernels by $F$ within $\mathsf{Ban}$, thus staying in the functional-analytic context.
@@
For a Banach $\Gamma$-module $V$, an almost invariant sequence is a sequence of vectors $v_{n}\in V$ such that for every $g\in\Gamma$, $\lim_{n}(1-g)v_{n}=0$. We end this subsection with the following observation, see [15]*Section 3.